OpenStudy (anonymous):

Bacteria A has an initial population of 500 and doubles every day, while bacteria b has an initial population of 50 and triples daily. A) After how long will the population of B overtake the population of A?

OpenStudy (anonymous):

so let a(t) be the Bacteria A and let b(t) be the Bacteria B so we have a(t)=500*2^t b(t)=50*3^t so we want to know when b(t)>a(t) so we get 50*3^t-500*2^t>0 divide by 50 3^t-10*2^t>0 3^t>10*2^t divide by 2^t (3/2)^t>10 use logarithm with base 3/2 t>lg(3/2,10) so in terms of natural logarithm you have t > ln(10)/ln(3/2)

OpenStudy (anonymous):

which is equal to roughly 5.67 that is you can say after 6 days it will be

OpenStudy (anonymous):

Got it, thank you!